# HippoMM: Hippocampal-inspired Multimodal Memory for Long Audiovisual Event Understanding

### v1

@@ 5.1 Implementation Details
Our system, HippoMM, was implemented on a NVIDIA L40s GPU. For Memory Formation, adaptive temporal segmentation processed video frames at 1-10 fps, identifying boundaries using SSIM and audio energy thresholds to create 5-10s segments. Perceptual Encoding generated ShortTermMemory objects ($m_{i}$, Eq. 7) containing 1024-dim ImageBind [girdhar2023imagebind] embeddings ($\mathcal{E}_{c}$), Whisper [radford2023robust] transcriptions ($\mathcal{T}_{a}$), Qwen2.5-VL [Qwen2.5-VL] visual descriptions ($\mathcal{T}_{v}$), and raw data pointers ($\mathbf{C}_{i}$). Memory Consolidation filtered segments based on cosine similarity between average ImageBind embeddings ($\mathbf{v}_{i}$) using threshold $\gamma$. Semantic Replay employed Qwen2.5-VL as $\phi_{\text{LLM}}$ to generate textual summaries ($\mathbf{S}_{\theta_{k}}$) stored in persistent ThetaEvent objects ($\theta_{k}$). For Memory Retrieval, GPT-4o [hurst2024gpt] served as the query classifier ($\mathcal{Q}_{\text{type}}$), directing queries initially to Fast Retrieval ($\Phi_{\text{fast}}$) over ThetaEvent summaries. If confidence was below $\tau=0.75$, Detailed Recall ($\Psi_{\text{detailed}}$) was invoked, using ImageBind embeddings for feature search and GPT-4o for semantic search over descriptions ($\mathbf{T}_{k}$). GPT-4o was also used as $\Gamma$ for final answer synthesis. Additional hyperparameters ($\tau_{v}$, $\tau_{a}$, $k$, $\delta$) are detailed in the Appendix.
@@
Figure 3: Visualization of consolidated ThetaEvent embedding space and query retrieval via t-SNE projection. (Left) Embeddings for visual features, auditory features, captions, and transcriptions in one event. The semantic summary ($\mathbf{S}_{\theta}$, green star) is central. (Right) Text query embeddings (‘dorm’, ‘football’ - stars) retrieve closest caption/transcription embeddings (crosses), linking queries to specific multimodal segments (corresponding frames and text shown).
5.2 Main Results
We evaluated HippoMM on the HippoVlog benchmark (introduced in Section 3) designed for long-form audiovisual understanding, comparing it against NotebookLM [^2] https://notebooklm.google.com/ and Video RAG [ren2024videorag]. These baselines were selected as they represent the only existing systems capable of processing both long-form video and audio content. As summarized in Table 1, HippoMM achieves a state-of-the-art average accuracy of 78.2%, significantly outperforming the next-best method (Video RAG) by 14.0%. This advantage is particularly evident in visual tasks (66.8% vs. Video RAG’s 41.2%), which we attribute to our adaptive memory consolidation preserving salient details more effectively than standard chunking. Furthermore, HippoMM demonstrates strong cross-modal reasoning (70.8% A+V accuracy), confirming its ability to capture inter-sensory temporal relationships [teyler2007hippocampal]. Note that PT and ART metrics are not reported for NotebookLM, as it operates as a web-based service potentially using pre-indexed knowledge, making direct comparison infeasible. In terms of efficiency, HippoMM’s adaptive sampling reduces preprocessing time (PT) by 46.2% compared to Video RAG (5.09h vs 9.46h). Critically, HippoMM maintains a balanced average response time (ART) of 20.4s. In contrast, Video RAG exhibits a much higher ART (112.5s), largely because its retrieval mechanism fails to locate relevant segments in 55.4% of cases, defaulting to costly detailed processing across the entire long video. HippoMM’s hierarchical retrieval avoids this pitfall, balancing responsiveness and accuracy.
5.3 Ablation Studies
To assess the contribution of individual components, we performed ablation studies by selectively removing Detailed Recall (DR), Fast Retrieval (FR), and Adaptive Reasoning (AR), with results in Table 1 (bottom rows). Removing detailed, frame-level analysis (variant ‘w/o DR, AR’) achieved the fastest ART (4.14s) but at the cost of a >6% drop in visual accuracy (to 60.4%), underscoring DR’s role in fine-grained visual understanding. Conversely, removing fast, summary-based access (variant ‘w/o FR, AR’) slightly improved cross-modal accuracy (to 72.0%) but drastically increased ART to 27.3s, confirming FR’s effectiveness for rapid gist retrieval. Finally, removing only the adaptive reasoning module (‘w/o AR’) resulted in a noticeable drop in overall accuracy (to 76.8%) and semantic understanding (-1.6%), highlighting AR’s function in refining and reconciling information from different retrieval paths. These ablations demonstrate that the synergy between rapid summary access (FR), detailed segment analysis (DR), and intelligent result integration (AR) is crucial for HippoMM’s robust performance. The experimental results collectively validate our central hypothesis: translating principles inspired by human memory yields a computational framework with significant advantages in accuracy and efficiency for complex, long-form audiovisual understanding tasks.
5.4 Analysis of HippoMM Mechanisms
5.4.1 Visualizing Consolidated Event Representations and Retrieval
HippoMM employs memory consolidation and semantic replay to transform sequences of short-term memories into coherent, semantically rich ThetaEvent representations. Figure 3 provides a visualization of one such ThetaEvent, projected into a 2D space using t-SNE based on ImageBind embeddings. The left panel shows embeddings from different constituents of the event: visual features (purple squares), auditory features (orange diamonds), derived captions (blue markers), and audio transcriptions (pink circles). Notably, the embedding of the generated summary ($\mathbf{S}_{\theta}$, green star) occupies a central position relative to the various modality-specific feature clusters, suggesting its effectiveness in integrating and balancing information across different perceptual inputs. The right panel demonstrates the utility of this representation for query-driven retrieval. Text queries, such as “dorm” (yellow star query embedding) and “football” (purple star query embedding), are projected into the same space. The system identifies the nearest neighbor embeddings within the detailed ShortTermMemory features (marked by yellow and purple crosses respectively). For the “dorm” query, the system successfully retrieves a semantically relevant transcription about “stay” which, while not explicitly containing the word “dorm,” correctly identifies the residential setting depicted in the corresponding visual frame. For the “football” query, it precisely retrieves a transcription explicitly mentioning the “first football game” and correctly identifies the associated football field visual scene. This illustrates how the consolidated ThetaEvent structure supports both abstract summarization and effective retrieval of specific, grounded details from the original multimodal stream.
5.4.2 Hierarchical Retrieval Dynamics
Our hierarchical retrieval (Section 4.3) aims to balance speed and accuracy by first attempting fast, summary-based retrieval ($\Phi_{\text{fast}}$) before resorting to detailed segment search ($\Psi_{\text{detailed}}$). We analyzed the usage and performance of these pathways across the HippoVlog benchmark queries, visualizing the results in Figure 4. This figure displays individual query outcomes plotted against their response times, differentiating pathways by color and correctness by marker style (filled for correct, open for incorrect). The analysis revealed that fast retrieval successfully handled 69.1% of queries (visible by the density of points below the 30s threshold), demonstrating the effectiveness of the semantic summaries in ThetaEvent objects for rapid information access. Detailed retrieval was primarily invoked for queries requiring fine-grained perceptual information, constituting the remaining 30.9% of cases (points at or above 30s). The dashed line marks the 30s threshold. As shown in the figure annotations, the two pathways exhibit significant differences in processing efficiency, with fast retrieval achieving an average response time of 6.5s compared to detailed retrieval’s 51.6s. This clear bimodal distribution validates our hierarchical design, confirming the system effectively routes queries to the appropriate pathway based on information needs.
Figure 4: Retrieval pathway analysis with x-axis showing response time. Blue markers ($<$30s) and purple markers ($\geq$30s) represent fast and detailed pathways respectively; filled markers indicate correct outcomes, open markers incorrect. Dashed line marks the 30s threshold.
5.5 Generalization Across Audiovisual Tasks
To assess HippoMM’s generalization capabilities and benchmark its performance against contemporary methods, we established a multifaceted evaluation protocol as detailed in Table 2. This protocol integrates metrics from two complementary domains to thoroughly evaluate multimodal understanding. First, we assess event representation fidelity through Content Relevance, Semantic Similarity, and Action Alignment metrics, leveraging ground-truth action scripts from the Charades dataset [sigurdsson2016hollywood] (a detailed collection of annotated home activity videos) with evaluations supported by GPT-4o [hurst2024gpt]. Second, we probe the long-duration temporal understanding of models via the Needle QA accuracy metric, derived from the challenging Needle-in-a-Haystack task in the MLVU benchmark [MLVU], which involves locating precise information within lengthy video sequences ($\geq$ 5 min) through 415 targeted queries. The comparison includes leading multimodal systems such as Qwen2.5-Omni [xu2025qwen2] and VideoLlama2 [damonlpsg2024videollama2], alongside GPT-4o serving as a ceiling reference. Evaluation constraints include GPT-4o processing only subsampled frames (0.5 fps), and Qwen2.5-Omni’s inability to complete the Needle QA task due to memory constraints.
Results clearly highlight HippoMM’s strong generalization capabilities. Individual metrics are reported along with a consolidated Overall Score, computed as the arithmetic mean after linearly normalizing the original 1–5 scales to a 0–100% range (note that Qwen2.5-Omni’s score excludes Needle QA). As presented in Table 2, HippoMM achieves the highest Overall Score (66.9%) among directly comparable systems. Crucially, HippoMM demonstrates balanced strengths, excelling in detailed visual metrics such as Action Alignment (top score of 3.14), while significantly outperforming all other evaluated methods on the demanding long-video Needle QA task (73.1%). Moreover, HippoMM closely approximates GPT-4o’s ceiling performance, underscoring its robust visual comprehension. Collectively, these findings emphasize HippoMM’s exceptional ability to generalize across complex and diverse multimodal contexts.
@@ 5.5 Generalization Across Audiovisual Tasks
Table 2: Comparison of mean evaluation scores. R: Relevance, S: Similarity, A: Alignment (Scores 1-5). NQA: Needle QA (%), Overall: Overall Score (%). GPT-4o is included as a ceiling reference for visual understanding performance. Higher values indicate better performance. Best results among directly comparable systems are in bold, second best are underlined.
@@ 6 Conclusion
We introduced HippoMM, a hippocampus-inspired architecture that translates neurobiological principles into computational mechanisms for multimodal memory. Our short-to-long term memory formation process enables efficient consolidation of perceptual details into semantic abstractions. While HippoMM outperforms existing methods on audiovisual benchmarks, it struggles with ambiguous event boundaries and complex cross-modal associations under noisy conditions. The dual-pathway retrieval system effectively balances efficiency and detail but requires further refinement to handle truly open-domain queries. This work demonstrates both the promise and limitations of neuroinspired approaches, revealing that despite progress in biomimetic design, substantial gaps remain between computational implementations and the robust, flexible memory capabilities observed in biological systems.
@@
Appendix A HippoVlog Construction Details
@@
A.1 Video Selection and Preprocessing
  • Source Selection: 25 daily vlogs from YouTube (Creative Commons licensed), with average duration of 15.3 minutes, totaling approximately 682 minutes of content.
@@ A.1 Video Selection and Preprocessing
A.2 Multimodal Data Sampling
@@
A.3 Segment-Level Ground Truth Generation
@@ A.4 Question Generation
A.5 Question Curation and Finalization
@@
A.6 Question Categories and Examples
@@ A.6.1 Cross-Modal Binding (TV×$\times$×A)
Example: Cross-Modal Binding Question: “When the main character is sitting in the car wearing a black jacket and white scarf, what does she audibly express?” Options: A: She says she feels excited about the drive. B: She mentions being scared and says it feels weird. C: She talks about the weather being too cold. D: She expresses confidence about her driving skills. Explanation: In the video, as the main character sits in the car wearing a black jacket and white scarf, she audibly expresses her nervousness by saying ‘okay, hey, I’m scared, oh my gosh this feels weird.’
A.6.2 Auditory-Focused Retrieval (TA)
@@
Example: Auditory-Focused Retrieval Question: “What did the main character mention about the purpose of the meeting?” Options: A: To finalize a business deal B: To discuss the upcoming project C: To plan a surprise party D: To talk about a vacation Explanation: The main character specifically mentions the ‘upcoming project’ as the purpose of the meeting, indicating a discussion centered around future plans and tasks.
A.6.3 Visual-Focused Retrieval (TV)
@@
Example: Visual-Focused Retrieval Question: “What distinctive pattern is visible on the roof of the building?” Options: A: Checkerboard B: Chevron C: Striped D: Polka dot Explanation: The roof of the building in the image has a distinctive chevron pattern.
Example: Visual-Focused Retrieval Question: “What is the main character wearing on her head?” Options: A: A baseball cap B: A knitted hat C: A colorful headscarf D: A headband Explanation: The main character is wearing a colorful headscarf, as seen in the image.
A.6.4 Semantic/Temporal Reasoning (TS)

### v2

@@ 5.1 Implementation Details
Our system, HippoMM, was implemented on an NVIDIA L40S GPU. Episodic segmentation processed video frames at 1-10 fps using SSIM and audio energy thresholds to create 5-10s segments. Perceptual Encoding generated ShortTermMemory objects ($m_{i}$, Eq. 4) containing 1024-dim ImageBind embeddings [9] ($\mathcal{E}_{c}$), Whisper transcriptions [27] ($\mathcal{T}_{a}$), Qwen2.5-VL visual descriptions [2] ($\mathcal{T}_{v}$), and raw data pointers ($\mathbf{C}_{i}$). Memory Consolidation filtered segments based on cosine similarity ($\gamma=0.85$) between average ImageBind embeddings. Qwen2.5-VL [2] generated textual summaries ($\mathbf{S}_{\theta_{k}}$) for ThetaEvent objects. For Hierarchical Memory Retrieval, GPT-4o [15] served as query classifier and final answer synthesizer ($\Gamma$), while Qwen2.5-VL provided confidence scoring during Fast Retrieval ($\tau=0.75$ threshold) and fine-grained visual analysis during Detailed Recall. For complete configurations and prompt templates, see Appendix.
@@
Figure 3: Visualization of consolidated ThetaEvent embedding space and query retrieval via t-SNE projection. (Left) Embeddings for visual features, auditory features, captions, and transcriptions in one event. The semantic summary ($\mathbf{S}_{\theta}$, green star) is central. (Right) Text query embeddings (“dorm,” “football” - stars) retrieve closest caption/transcription embeddings (crosses), linking queries to specific multimodal segments (corresponding frames and text shown).
5.2 Main Results and Ablation Studies
HippoMM yields superior accuracy and efficiency. Table 1 presents our results on the HippoVlog benchmark, designed to evaluate long-form, cross-modal associative recall. HippoMM achieves state-of-the-art 78.2% accuracy, outperforming strong baselines like Ola [21] (74.9%) and VideoLLaMA 3 [41] (73.0%). Notably, we implemented a parameter-free cognitive baseline, Simple-Assoc-Memory (SAM), which uses classic Hebbian auto-association on the same episodic features as HippoMM. Its poor performance (30.3%) confirms that a naive cognitive mapping is insufficient, validating the necessity of our structured architectural components. Furthermore, HippoMM demonstrates superior efficiency: 46.2% reduction in processing time and $5\times$ faster response time than VideoRAG [28], validating that biological structure enables both accuracy and efficiency gains.
Ablation studies in Table 1 validate each component. Removing the Detailed Recall path (HippoMM w/o DR) causes accuracy to drop to 61.2%, confirming its necessity for fine-grained perceptual reasoning. Conversely, removing the Fast Retrieval path (HippoMM w/o FR) maintains accuracy (74.6%) but increases average response time to 19.54s, demonstrating the semantic index’s critical role in efficiency. A pure Embedding-Only Retrieval (EOR-only) baseline still achieves 71.1%, proving the inherent strength of the underlying memory structure independent of LLM reasoning. Finally, replacing GPT-4o with the smaller Qwen2.5-14B maintains strong performance (70.8%), confirming that our cognitive architecture rather than model scale drives effectiveness.
5.3 Evaluation of Generalization and Core Capabilities
Episodic segmentation enables better temporal understanding. We evaluated HippoMM through a comprehensive protocol spanning event representation and temporal understanding (Table 2). For event representation, we measure Content Relevance (R), Semantic Similarity (S), and Action Alignment (A) using Charades [30] with GPT-4o-based evaluation. For temporal understanding, we employ the Needle-in-a-Haystack (NQA) task from MLVU [43], which requires locating specific information within videos exceeding five minutes. HippoMM achieves 73.1% on NQA, a 46% improvement over VideoLlama 2, while also leading in Semantic Similarity and Action Alignment. Qwen2.5-Omni performs well on shorter videos but did not yield NQA results under our evaluation setup. These results validate that hippocampus-inspired episodic segmentation enables superior temporal navigation. On Video-MME [7] (61.8%) and LongVideoBench [36] (55.2%), HippoMM shows competitive performance; full cross-benchmark comparison is in the Appendix (Table 4).
@@ 5.3 Evaluation of Generalization and Core Capabilities
Table 2: Evaluation scores across core capabilities. R: Relevance, S: Similarity, A: Alignment (1-5 scale). NQA: Needle QA (%). Overall: Mean normalized score (%). GPT-4o included as a strong reference baseline. Best results in bold, second best underlined.
@@
5.4 Analysis of HippoMM Mechanisms
Consolidation enables cross-modal pattern completion. Figure 3 visualizes how memory consolidation creates unified ThetaEvent representations. The semantic summary ($\mathbf{S}_{\theta}$) occupies a central position among modality-specific clusters, integrating information across perceptual inputs. In retrieval, text queries like “dorm” successfully retrieve semantically relevant content (“stay” transcription with residential visuals) despite lexical mismatch, while “football” precisely retrieves “first football game” mentions with field visuals. This demonstrates how consolidated structures enable both abstract summarization and specific detail retrieval from partial cross-modal cues.
Hierarchical Memory Retrieval reduces computational demands. Figure 4 analyzes our hierarchical memory retrieval dynamics. Fast retrieval successfully handles 69.1% of queries with 6.5s average response time, while detailed retrieval processes complex queries requiring fine-grained perceptual information in 51.6s. This 8x speedup for the majority of queries, achieved without sacrificing accuracy, confirms that biological hierarchical structure yields computational efficiency through intelligent query routing. Our GPT-4o query classifier achieves 94.1% routing accuracy; even misrouted queries achieve 52.5% accuracy, demonstrating graceful degradation.
Figure 4: Retrieval pathway analysis with x-axis showing response time. Blue markers (<<30s) and purple markers ($\geq$30s) represent fast and detailed pathways respectively; filled markers indicate correct outcomes, open markers incorrect. Dashed line marks the 30s threshold.
Episodic segmentation preserves event integrity. We quantify segmentation quality through Event Integrity Rate (EIR), measuring the fraction of ground-truth events fully contained within single segments:
$\text{EIR}=\frac{1}{|E_{gt}|}\sum_{e\in E_{gt}}\mathbb{I}(\exists s\in S_{gen}:s\text{ fully contains }e)$ (13)
where $E_{gt}$ denotes ground-truth events, $S_{gen}$ denotes our generated segments, and $\mathbb{I}(\cdot)$ is the indicator function. On Charades [30] with ground-truth action annotations, our episodic segmentation achieves 61.6% EIR versus 40.2% for fixed 15-second chunks. This improvement demonstrates superior preservation of semantic event structure, crucial for memory formation and query accuracy.
Consolidation provides noise resilience. We evaluated HippoMM’s robustness by introducing varying levels of perceptual noise to the HippoVlog dataset. For visual noise, we applied ffmpeg’s noise filter with alls parameters of 10, 25, and 50 for low, medium, and high levels respectively. For auditory noise, we mixed white noise at 0.1, 0.3, and 0.5 volume ratios. Under visual noise, overall accuracy decreased from the clean baseline of 78.2% to 75.7% (low), 75.0% (medium), and 74.4% (high); under auditory noise, accuracy declined to 77.5%, 77.0%, and 74.7% respectively. Even under severe conditions, maximum degradation remains below 4 percentage points. This graceful degradation validates that consolidated multimodal representations provide effective redundancy against perceptual corruption, leveraging both pre-trained encoder robustness and the error-correcting properties of cross-modal consolidation.
Hyperparameter Robustness. We validated that our key hyperparameters ($\tau$ and $\tau_{v}$) are empirically optimal. The retrieval threshold ($\tau=0.75$) provides a clear speed-accuracy trade-off: lower values needlessly increase ART, while higher values cause accuracy to drop as detailed recall is bypassed. Similarly, the visual segmentation threshold ($\tau_{v}=0.65$) is at an optimal point: lower values over-segment and shatter events, while higher values under-segment and incorrectly merge them, with both cases degrading final accuracy. Our default values are thus robust and empirically justified.
@@ 6 Conclusion
HippoMM implements hippocampal principles of pattern separation, associative recall, and memory consolidation for multimodal understanding, achieving 78.2% accuracy on HippoVlog while maintaining 5× faster response than retrieval-augmented baselines. Ablations confirm each component’s necessity, while hierarchical retrieval balances efficiency with quality. This work demonstrates that cognitive architectures grounded in neuroscience provide principled foundations for multimodal computing systems, suggesting biological memory mechanisms offer promising pathways toward human-like understanding in artificial systems.
Acknowledgments
This work was supported in part by NSF CNS-2112562 and ARO W911NF-23-2-0224.
@@ Acknowledgments
Appendix A Key Notation
Table 3: Key notation used throughout the paper
Symbol Description
$X=\{x_{t}\}$ Continuous multimodal data stream
$\mathcal{S}_{t}$ Segmentation trigger indicator at time $t$
$\text{1}[\cdot]$ Indicator function (returns 1 if true, 0 otherwise)
$m_{i}$ $i$-th ShortTermMemory object
$K$ Set of indices for consolidated memories
$k\in K$ Index in the consolidated set
$m_{k}$ Consolidated ShortTermMemory object
$\theta_{k}$ $k$-th ThetaEvent object
$\mathcal{M}$ Collection of consolidated ShortTermMemory objects
$\Theta$ Collection of ThetaEvent objects
$\mathbf{v}_{i},\mathbf{v}_{k}$ Representative embedding for segment $i$ or $k$
$\mathbf{E}_{i}$ Sequence of cross-modal embeddings for segment $i$
$\mathbf{T}_{i}$ Aggregated textual outputs for segment $i$
$\mathbf{C}_{i}$ Pointers to raw audiovisual data for segment $i$
$\mathbf{S}_{\theta_{k}}$ Semantic summary of ThetaEvent $k$
$\mathcal{E}_{c}$ Cross-modal embedding model
$\mathcal{T}_{v}$ Vision-language model for visual descriptions
$\mathcal{T}_{a}$ Speech recognition model
$\phi_{\text{LLM}}$ LLM for semantic summary generation
$\gamma$ Consolidation similarity threshold
$\tau$ Confidence threshold for retrieval switching
$\delta$ Temporal window expansion parameter
$\cos(\cdot,\cdot)$ Cosine similarity function
$\text{sim}(\cdot,\cdot)$ Similarity function (cosine similarity)
$\text{TopK}(\cdot,k)$ Function returning $k$ highest-scoring elements
$\text{FormattedContext}(\cdot)$ Function to structure multimodal content
$\text{Extract}_{\text{target}}(\cdot)$ Function to extract target modality content
$(t_{s,i},t_{e,i})$ Start and end times of segment $i$
$t_{s,i},t_{e,i}]$ Alternative notation for segment boundaries
$\Phi_{\text{fast}}$ Fast retrieval function
$\Psi_{\text{detailed}}$ Detailed recall function
$\Gamma$ Adaptive reasoning module
$\mathcal{C}_{\text{context}}$ Context information for reasoning
Appendix B HippoVlog Construction Details
@@
B.1 Video Selection and Preprocessing
  • Source Selection: 25 daily vlogs from YouTube (Creative Commons licensed), with average duration of 27.3 minutes, totaling approximately 682 minutes of content.
@@ B.1 Video Selection and Preprocessing
B.2 Multimodal Data Sampling
@@
B.3 Segment-Level Ground Truth Generation
@@ B.4 Question Generation
B.5 Question Curation and Finalization
@@
B.6 Question Categories and Examples
@@
Example: Cross-Modal Binding
Question: “When the main character is sitting in the car wearing a black jacket and white scarf, what does she audibly express?”
Options:
  • A: She says she feels excited about the drive.
  • B: She mentions being scared and says it feels weird.
  • C: She talks about the weather being too cold.
  • D: She expresses confidence about her driving skills.
Explanation: In the video, as the main character sits in the car wearing a black jacket and white scarf, she audibly expresses her nervousness by saying ‘okay, hey, I’m scared, oh my gosh this feels weird.’
B.6.2 Auditory-Focused Retrieval (TA)
@@
Example: Auditory-Focused Retrieval
Question: “What did the main character mention about the purpose of the meeting?”
Options:
  • A: To finalize a business deal
  • B: To discuss the upcoming project
  • C: To plan a surprise party
  • D: To talk about a vacation
Explanation: The main character specifically mentions the ‘upcoming project’ as the purpose of the meeting, indicating a discussion centered around future plans and tasks.
B.6.3 Visual-Focused Retrieval (TV)
@@
Example: Visual-Focused Retrieval
Question: “What distinctive pattern is visible on the roof of the building?”
Options:
  • A: Checkerboard
  • B: Chevron
  • C: Striped
  • D: Polka dot
Explanation: The roof of the building in the image has a distinctive chevron pattern.
Example: Visual-Focused Retrieval
Question: “What is the main character wearing on her head?”
Options:
  • A: A baseball cap
  • B: A knitted hat
  • C: A colorful headscarf
  • D: A headband
Explanation: The main character is wearing a colorful headscarf, as seen in the image.
B.6.4 Semantic/Temporal Reasoning (TS)